OpenStudy (anonymous):

if (x-c) is a factor of 6x^3+3cx^2-c^2x-27, what is the value of c?

myininaya (myininaya):

if x-c is a factor then f(c)=0

myininaya (myininaya):

so whatever you have there as a function replace all the x's with c's

myininaya (myininaya):

and then set that expression equal to 0 and solve

OpenStudy (anonymous):

thank you! myininaya

myininaya (myininaya):

np you can show me what you did to get c and i can let you know if it is good or find your error or whatever

myininaya (myininaya):

i will be on for like a few more minutes

OpenStudy (anonymous):

I'm not sure, this is what i got. cube root of 27/8?

myininaya (myininaya):

yeah good and 27=3^3 and 8=2^3

myininaya (myininaya):

so you can simplify your answer

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

i don't get how to simplify

myininaya (myininaya):

\[\sqrt[3]{\frac{27}{8}}=\frac{\sqrt[3]{27}}{\sqrt[3]{8}}=\frac{\sqrt[3]{3^3}}{\sqrt[3]{2^3}}=\frac{3}{2}\]

myininaya (myininaya):

cube root cancels out the cube

OpenStudy (anonymous):

You're so smart! :)